# Towards Accurate and Interpretable Neuroblastoma Diagnosis via Contrastive Multi-scale Pathological Image Analysis

### v1

@@ Abstract
Neuroblastoma, adrenal-derived, is among the most common pediatric solid malignancies, characterized by significant clinical heterogeneity. Timely and accurate pathological diagnosis from hematoxylin and eosin-stained whole slide images is critical for patient prognosis. However, current diagnostic practices primarily rely on subjective manual examination by pathologists, leading to inconsistent accuracy. Existing automated whole slide image classification methods encounter challenges such as poor interpretability, limited feature extraction capabilities, and high computational costs, restricting their practical clinical deployment. To overcome these limitations, we propose CMSwinKAN, a contrastive-learning-based multi-scale feature fusion model tailored for pathological image classification, which enhances the Swin Transformer architecture by integrating a Kernel Activation Network within its multilayer perceptron and classification head modules, significantly improving both interpretability and accuracy. By fusing multi-scale features and leveraging contrastive learning strategies, CMSwinKAN mimics clinicians’ comprehensive approach, effectively capturing global and local tissue characteristics. Additionally, we introduce a heuristic soft voting mechanism guided by clinical insights to seamlessly bridge patch-level predictions to whole slide image-level classifications. We validate CMSwinKAN on the PpNTs dataset, which was collaboratively established with our partner hospital and the publicly accessible BreakHis dataset. Results demonstrate that CMSwinKAN performs better than existing state-of-the-art pathology-specific models pre-trained on large datasets. Our source code is available at https://github.com/JSLiam94/CMSwinKAN.
@@ 1 Introduction
Peripheral neuroblastic tumors (pNTs) are malignant tumors originating in the developing sympathetic nervous system and represent the most common solid extracranial tumors in children [1], accounting for 15% of all childhood cancer-related deaths [2, 3]. Due to the multiple subtypes of pNTs, significant differences exist in each subtype’s pathological characteristics and corresponding treatment strategies. Thus, precise subtype classification is critical for risk stratification and the development of personalized treatment plans. At present, pathological diagnosis is the gold standard for diagnosing pNTs. However, due to the subtle pathological differences between subtypes, diagnosis in clinical practice primarily relies on manual examination by pathologists, which is characterized by significant subjectivity and limitations [4, 5]. It is challenging to perform a holistic evaluation of whole slide images (WSIs). Therefore, developing automated methods to precisely identify cytopathological morphologies across entire WSIs and achieve accurate subtype diagnosis has become an urgent need.
@@
Previous studies have developed automated WSI diagnostic methods [6, 7], yet critical challenges persist. First, the feature representation capabilities of current methods are still insufficient. Conventional machine learning approaches remain limited by handcrafted features that cannot fully characterize pathological details and tissue variations. This fundamentally constrains model performance. Second, while deep learning methods have partially overcome the limitations of handcrafted feature extraction [8], their nonlinear modeling capabilities remain constrained. Specifically, when dealing with the highly heterogeneous and complex image data of histopathology, conventional convolutional neural network (CNN) architectures struggle to efficiently capture long-range spatial dependencies and multi-scale features in the images [9]. This may further hinder the robustness of these models in practical clinical applications [10]. Moreover, the lack of model interpretability is another critical factor impeding the clinical deployment of current models [11]. Deep learning models are often considered “black boxes” and lack transparent decision-making processes. This opacity increases clinical risks and erodes pathologists’ confidence in model outputs. Such interpretability limitations significantly hinder the adoption of automated diagnostic tools in clinical practice [12].
Additionally, the size of WSI images is immense, with a single image often reaching giga pixels, far exceeding the memory capacity of a graphics processing unit (GPU) [13, 14]. As a result, the current primary solution is to divide an entire WSI image into multiple independent patches for classification. However, this patch-level processing approach poses severe challenges related to information fragmentation and false associations [15, 16]. On the one hand, the lack of practical spatial constraints between patches may result in classification outcomes that overlook the overall spatial structure of the image. On the other hand, certain critical lesion regions may be divided into separate patches and processed in isolation, impairing the comprehension of the overall pathology and leading to incomplete or inaccurate classification. These factors significantly reduce the overall diagnostic accuracy and generalization ability of the models. Therefore, there is an urgent need for a WSI classification model that can better integrate local details and global contextual relationships, possess stronger feature representation and nonlinear modeling capabilities, and offer high interpretability to facilitate the practical clinical application of automated pathology diagnostic technologies.
@@
To address these challenges, this study proposes a novel contrastive-learning-driven multi-scale feature fusion pathological image classification model, CMSwinKAN, for classifying hematoxylin and eosin-stained(H&E) neuroblastoma (NB) histopathology patches. Additionally, a soft voting mechanism based on prior knowledge of tissue composition analysis is introduced to achieve accurate WSI classification.
Our main contributions are as follows:
  1. 1.
@@
  2. 2.
    Lightweight and interpretable deep architecture: We introduce a multi-head self-attention module with a combined configuration of regular and shifted windows (Swin KANsformer Block) and integrate a Kernel Activation Network (KAN) as the classification head. This significantly enhances the model’s feature representation and nonlinear modeling performance. Additionally, we optimize model parameters based on feature analysis, selecting appropriate embedding dimensions and the number of neurons in KAN’s hidden layers. This achieves a lightweight design while maintaining fine-grained classification accuracy.
  3. 3.
    Soft voting mechanism for WSI-level classification: We propose a WSI classification method based on clinical prior knowledge and a soft voting mechanism. By training a support vector machine (SVM) model, we achieve fine-grained recognition of tissue components at the patch level. Combined with a dynamic weight allocation strategy, this method effectively integrates local patch features with global features, improving classification accuracy and robustness.
@@ 2.1 Traditional Computer Vision Models in Pathology
Traditional approaches for pathological image classification depend on handcrafted feature engineering, struggling to extract discriminative patterns from high-resolution images. For example, Gheisari et al. [17]proposed a method combining Scale-Invariant Feature Transform (SIFT) with feature encoding strategies, using an SVM classifier to classify NB pathological images into five subtypes. However, its performance was limited by the expressiveness of handcrafted features, achieving an accuracy of only 68.2%. A subsequent improvement introduced a Convolutional Deep Belief Network (CDBN) for NB pathological image classification. The CDBN model extracted high-level patterns via hierarchical feature encoding, achieving an F1 score of 86.01% under the Shimada classification system [18]. However, due to the extremely high resolution of WSIs, this approach faced challenges such as low computational efficiency, high resource requirements for whole-slide processing, and the inability to effectively model long-range spatial dependencies.
@@ 2.2 Development of Pathology Large Models
Recent advances in self-supervised learning (SSL) have propelled breakthroughs in large pathology models. For instance, Wang et al. proposed the CTransPath model, which combines CNNs with a multi-scale Swin Transformer architecture, achieving state-of-the-art (SOTA) performance across multiple pathology tasks [21]. Additionally, Chen et al. introduced the universal pathology pretraining model UNI, which was pre-trained on over 1 million pathological images. UNI surpassed existing methods across 34 pathology tasks, achieving comprehensive breakthroughs from tissue classification to cancer subtype prediction [22].
The introduction of vision-language models has further enhanced the generalization capabilities of large pathology models. Lu et al. proposed the CONCH model, which was pre-trained on 1.17 million image-text pairs in a task-agnostic manner, achieving leading performance in tasks such as pathological image classification, segmentation, and retrieval [23]. Furthermore, Azizi et al. developed the REMEDIS framework, which improved the robustness of medical image models through SSL, enabling high performance even in out-of-distribution scenarios [24].
@@ 2.3 Advantages of Lightweight Small Models
Despite strong performance, pathology large models face four key limitations: high computational resource consumption (e.g., months of GPU pretraining), limited deployability on medical edge devices, lack of interpretability (due to complex model structures that are difficult to map to clinically interpretable pathological features), and poor adaptability to small-sample datasets, such as pediatric tumors.
In contrast, lightweight, small models offer unique advantages in terms of computational efficiency and adaptability. For example, Zormpas-Petridis et al. proposed the SuperHistopath framework, which uses the Simple Linear Iterative Clustering superpixel algorithm to segment low-resolution WSIs and classify superpixels based on CNNs, enabling precise mapping of tumor heterogeneity [25]. Ding et al. introduced the Fuzzy-Guided Multi-Granularity Deep Neural Network (FMDNN), which simulates the multi-level diagnostic process of pathologists and outperformed existing classification methods on several public datasets [26].
In the field of efficient Transformer design, Stegmüller et al. proposed ScoreNet, which uses a differentiable recommendation mechanism to focus on the most discriminative pathological regions and incorporates the ScoreMix data augmentation strategy to enhance generalization [27]. Additionally, Liu et al. developed the DetexNet network, which combines Texture Energy Measure to improve the representation ability of low-level pathological features, enabling high accuracy on small-sample pathology datasets [28]. Although these methods are significantly more efficient than large models, they still lack sufficient capacity to model local features such as nuclear details and tissue structure patterns while effectively integrating multi-scale information.
@@ 2.4 Multiscale Methods in Medical Image Processing
Multiscale feature extraction and fusion have long been critical for improving the performance of image processing models. In the field of natural images, Chen et al. proposed the CrossViT model, which uses a dual-branch Transformer to process image patches of different sizes and fuses them via a cross-attention mechanism, effectively improving image classification [29]. Li et al. introduced MViTv2, which improved multiscale visual Transformers by decomposing relative position encoding and incorporating residual pooling connections, achieving efficient performance in classification and detection tasks [30]. Ren et al. proposed Shunted Self-Attention, which includes heterogeneous receptive fields at each layer to aggregate multiscale features of tokens, effectively addressing the limitations of traditional self-attention in handling multiscale objects [31]. Additionally, Jiao et al. introduced DilateFormer, which uses multiscale dilated attention to expand the receptive field while maintaining computational efficiency, achieving SOTA performance across multiple vision tasks [32].
In the medical image domain, Huo et al. proposed the Hierarchical Multiscale Feature Fusion Network (HiFuse), which integrates global and local feature blocks and uses adaptive feature fusion modules to combine information from different scales efficiently, significantly enhancing model robustness in complex medical image classification tasks [33]. Chen et al. proposed an accurate leukocyte detection model based on deformable-DETR integrated with multi-level feature fusion, effectively capturing detailed cellular characteristics through deformable self-attention and hierarchical fusion mechanisms, thus significantly enhancing leukocyte detection accuracy [34]. Lei et al. introduced the ConDSeg framework, which uses a contrast-driven feature enhancement strategy to improve boundary recognition in image segmentation tasks, achieving SOTA performance on multiple pathology datasets [35]. Given its outstanding performance in medical image segmentation, we incorporate similar strategies into pediatric pathology image classification to address challenges such as multiscale information fusion and boundary ambiguity, thereby improving classification accuracy and model generalization capabilities.
@@ 3.1 Overall Architecture
Fig. 1 illustrates the overall architecture of the CMSwinKAN model (micro version), which is designed to address the challenges of multi-scale feature representation and spatial dependency modeling in pathological image classification. This model is specifically tailored for the patch-level classification of pathological images from H&E-stained slides by combining a lightweight architecture with clinical interpretability. The model consists of three core modules:
1). Swin KANsformer Block: This is an improved computational unit based on the Swin Transformer, where the traditional multi-layer perceptron (MLP) layers are replaced with KAN layers. This substitution effectively boosts both local feature representation and interpretability.
2). CMSA Module: By leveraging contrastive learning and an adaptive feature weighting mechanism, this module integrates multi-scale pathological patterns (e.g., nuclear details and tissue structures), thereby simulating the comprehensive diagnostic strategies of clinical pathologists.
3). KAN Classification Head: The KAN layers replace the conventional fully connected layers in the classification head, improving the model’s non-linear representation capacity while reducing computational complexity.
With the introduction of KAN layers, the model significantly reduces parameter count by utilizing smaller embedding dimensions (e.g., 24 for the micro version and 56 for the tiny version, compared to 128 for Swin-base). This reduction comes with minimal performance degradation; in fact, the feature representation and generalization capabilities are enhanced. Furthermore, the Swin KANsformer Block alternates between Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA) layers, replacing all MLP layers with KAN layers, which further improves the model’s efficiency and interpretability.
The core structure of Contrastive-Driven Feature Augmentation (CDFA) in the CMSA module enhances multi-scale feature fusion by comparing attention across different scales, resulting in richer and more diagnostically meaningful insights. The three modules work synergistically, enabling the CMSwinKAN model to efficiently and accurately classify H&E-stained pathological image patches and laying the foundation for whole-slide image classification tasks.
@@ 3.2 Swin KANsformer Block
The Swin Transformer Block has demonstrated numerous advantages in image processing and medical image analysis compared to traditional Vision Transformers. Introducing hierarchical feature maps and a local window-based self-attention mechanism effectively reduces computational complexity while enabling the extraction of multi-scale features, making it particularly well-suited for handling high-resolution images. Additionally, its shifted window mechanism enhances information exchange between windows, improving the model’s representation and feature extraction capabilities. However, the Swin Transformer Block has limitations, such as constrained global receptive fields, high computational complexity, and reliance on large-scale pretraining datasets. To address these shortcomings, we present the Swin KANsformer Block, an optimized enhancement of the original Swin Transformer Block.
The Swin KANsformer Block replaces the traditional MLP layers in the Swin Transformer Block with KAN layers, addressing the deficiencies of the Swin Transformer when processing complex H&E-stained pathological image data. Inheriting the structure of the Swin Transformer Block, its multi-head self-attention layers alternate between two types in consecutive Swin KANsformer Blocks: W-MSA layers and SW-MSA layers. Within the model, the input RGB H&E-stained pathological images are first divided into non-overlapping 4×4 image patches via the Patch Partition module, with each patch having a feature dimension of 4×4×3 = 48. A linear embedding layer then projects these patches into an arbitrary feature dimension, denoted as C.
@@ 3.3 Contrastive-Driven Multi-Scale Aggregation Module
Analyzing images at multiple scales leverages scale-aware knowledge to enhance model performance and uses the learned cross-scale relationships to maintain spatial consistency. We employ the CMSA module to utilize multi-scale feature information effectively. This module introduces the CDFA module, which addresses two significant challenges in medical image analysis: “soft boundaries” and co-occurrence phenomena. In our model, the CDFA module serves as the fusion mechanism for multi-scale features extracted from different layers of the Swin KANsformer.
@@
$\hat{\bm{V}}_{\Delta_{i,j}}=Softmax(\hat{A}_{1_{i,j}})\otimes(Softmax(\hat{A}_{2_{i,j}})\otimes(\bm{V}_{\Delta_{i,j}}),$ (3)
@@ 3.4 KAN Classification Head
The multi-scale fused features obtained from the CMSA module are concatenated with the output residuals from the 4th stage of the Swin KANsformer. This step appropriately preserves the detail and semantic information from the original image, enabling the multi-scale fused features to guide image classification effectively. We employ the KAN as the classification head to endow the model with strong non-linear fitting capabilities and superior classification performance. The KAN processes input features and outputs classification probabilities. Each activation function in KAN is a learnable spline function, which can be expressed as:
@@
KAN uses B-spline basis functions to construct piecewise polynomials with local support (i.e., each basis function is non-zero only within a finite interval). This property allows the model to handle high-dimensional data efficiently.
The B-spline basis function serves as the core component of KAN for constructing piecewise polynomials. For input data $x$ and a given set of grid points $t=\{t_{0},t_{1},\ldots,t_{n}\}$, the B-spline basis function $N_{i,k}\left(x\right)$ is computed as follows:

### v2

@@ Abstract
Neuroblastoma, adrenal-derived, is among the most common pediatric solid malignancies, characterized by significant clinical heterogeneity. Timely and accurate pathological diagnosis from hematoxylin and eosin-stained whole-slide images is critical for patient prognosis. However, current diagnostic practices primarily rely on subjective manual examination by pathologists, leading to inconsistent accuracy. Existing automated whole-slide image classification methods encounter challenges such as poor interpretability, limited feature extraction capabilities, and high computational costs, restricting their practical clinical deployment. To overcome these limitations, we propose CMSwinKAN, a contrastive-learning-based multi-scale feature fusion model tailored for pathological image classification, which enhances the Swin Transformer architecture by integrating a Kernel Activation Network within its multilayer perceptron and classification head modules, significantly improving both interpretability and accuracy. By fusing multi-scale features and leveraging contrastive learning strategies, CMSwinKAN mimics clinicians’ comprehensive approach, effectively capturing global and local tissue characteristics. Additionally, we introduce a heuristic soft voting mechanism guided by clinical insights to seamlessly bridge patch-level predictions to whole-slide image-level classifications. We validate CMSwinKAN on the PpNTs dataset, which was collaboratively established with our partner hospital and the publicly accessible BreakHis dataset. Results demonstrate that CMSwinKAN performs better than existing state-of-the-art pathology-specific models pre-trained on large datasets. Our source code is available at https://github.com/JSLiam94/CMSwinKAN.
@@ 1 Introduction
Peripheral neuroblastic tumors (pNTs) are malignant tumors originating in the developing sympathetic nervous system and represent the most common solid extracranial tumors in children [1], accounting for 15% of all childhood cancer-related deaths [2]. Due to the multiple subtypes of pNTs, significant differences exist in each subtype’s pathological characteristics and corresponding treatment strategies. Thus, precise subtype classification is critical for risk stratification and the development of personalized treatment plans. At present, pathological diagnosis is the gold standard for diagnosing pNTs. However, due to the subtle pathological differences between subtypes, diagnosis in clinical practice primarily relies on manual examination by pathologists, which is characterized by significant subjectivity and limitations [3, 4]. It is challenging to perform a holistic evaluation of whole-slide images (WSIs). Therefore, developing automated methods to precisely identify cytopathological morphologies across entire WSIs and achieve accurate subtype diagnosis has become an urgent need.
@@
Previous studies have developed automated whole slide imaging (WSI) diagnostic methods [5, 6], but several challenges remain. First, current feature representation capabilities are inadequate, with conventional machine learning relying on handcrafted features that fall short in capturing pathological details and tissue variations. While deep learning has improved upon feature extraction [7], conventional convolutional neural networks (CNNs) still struggle with the complexity of histopathology images, limiting their ability to capture long-range dependencies and multi-scale features[8]. This issue can reduce the robustness of these models in clinical settings [9]. Furthermore, the lack of model interpretability complicates clinical deployment, as deep learning models often act as ”black boxes,” leading to increased risks and diminished pathologist confidence in their outputs [10]. These interpretability challenges significantly limit the use of automated diagnostic tools in practice [11].
The size of WSI can reach giga pixels, which often exceeds a graphics processing unit (GPU) memory capacity [12]. Consequently, the common practice is to divide these images into multiple patches for classification. However, this patch-level processing leads to challenges like information fragmentation and false associations [13]. It may overlook essential spatial structures and disrupt critical lesion regions, resulting in reduced diagnostic accuracy. Thus, a classification model with strong interpretability is urgently needed that better integrates local details with global context, offers strong feature representation for clinical applications.
@@
To tackle these issues, this study proposes CMSwinKAN, a contrastive multi-scale feature fusion model for patch-level classification in pNTs subtyping, with a tissue-aware soft voting mechanism for WSI-level analysis on hematoxylin and eosin (H&E)-stained pathological images. Our main contributions are as follows:
@@
  • Lightweight and interpretable deep architecture: We introduce a multi-head self-attention module with a combined configuration of regular and shifted windows (Swin KANsformer Block) and integrate a Kernel Activation Network (KAN) as the classification head. This significantly enhances the model’s feature representation and nonlinear modeling performance.
  • Soft voting mechanism for WSI-level classification: We propose a WSI classification method using clinical prior knowledge and a soft voting mechanism. By training a support vector machine (SVM), we achieve fine-grained tissue component recognition at the patch level. This approach integrates local and global features through a dynamic weight allocation strategy, enhancing classification accuracy and robustness.
@@ 2.1 Traditional Computer Vision Models in Pathology
Traditional methods for pathological image classification rely on handcrafted features, which struggle with high-resolution images. Gheisari et al. [14] used Scale-Invariant Feature Transform and SVM to classify NB pathological images into five subtypes but only achieved 68.2% accuracy due to limited feature expressiveness. An improvement with a Convolutional Deep Belief Network (CDBN) extracted higher-level patterns and reached an F1 score of 86.01% using the Shimada classification system [15]. However, this approach faced challenges like low computational efficiency and difficulty modeling long-range spatial dependencies due to the high resolution of WSIs.
@@ 2.2 Development of Pathology Large Models
Recent advances in self-supervised learning (SSL) have propelled breakthroughs in large pathology models. For instance, Wang et al. [18] proposed the CTransPath model, which combines CNNs with a multi-scale Swin Transformer architecture, achieving state-of-the-art (SOTA) performance across multiple pathology tasks. Additionally, Chen et al. introduced the universal pathology pretraining model UNI, which was pre-trained on over 1 million pathological images. UNI surpassed existing methods across 34 pathology tasks, achieving comprehensive breakthroughs from tissue classification to cancer subtype prediction [19].
The introduction of vision-language models has further enhanced the generalization capabilities of large pathology models. Lu et al. [20] proposed the CONCH model, which was pre-trained on 1.17 million image-text pairs in a task-agnostic manner, achieving leading performance in tasks such as pathological image classification, segmentation, and retrieval. Furthermore, Azizi et al. [21] developed the REMEDIS framework, which improved the robustness of medical image models through SSL.
@@ 2.3 Advantages of Lightweight Small Models
Pathology large models exhibit four main limitations: high computational resource demands (e.g., months of GPU pretraining), limited deployability on medical edge devices, lack of interpretability, and poor adaptability to small-sample datasets like pediatric tumors.
In contrast, lightweight, small models offer unique advantages in terms of computational efficiency and adaptability. For example, Zormpas-Petridis et al. [22] proposed the SuperHistopath framework, which uses the Simple Linear Iterative Clustering superpixel algorithm to segment low-resolution WSIs and classify superpixels based on CNNs, enabling precise mapping of tumor heterogeneity. Ding et al. [23] introduced the Fuzzy-Guided Multi-Granularity Deep Neural Network (FMDNN), which effectively simulates pathologists’ multi-level diagnostic processes and performs well on public datasets.
In the field of efficient Transformer design, Stegmüller et al. [24] proposed ScoreNet, which uses a differentiable recommendation mechanism to focus on the most discriminative pathological regions and incorporates the ScoreMix data augmentation strategy to enhance generalization. Additionally, Liu et al. [25] developed the DetexNet network, which combines Texture Energy Measure to improve the representation ability of low-level pathological features, enabling high accuracy on small-sample pathology datasets. These methods are more efficient than large models but still struggle to capture local features like nuclear details and tissue patterns while integrating multi-scale information.
@@ 2.4 Multiscale Methods in Medical Image Processing
Multiscale feature extraction and fusion are essential for enhancing image processing models. Chen et al. [26] proposed the CrossViT model, which uses a dual-branch Transformer with a cross-attention mechanism to fuse image patches of different sizes, improving image classification. Li et al. [27] introduced MViTv2, enhancing visual Transformers through relative position encoding and residual pooling connections。 Ren et al. [28] presented Shunted Self-Attention, which aggregates multiscale features with heterogeneous receptive fields, overcoming the limitations of traditional self-attention. Jiao et al. [29] developed DilateFormer, utilizing multiscale dilated attention for efficient performance across various vision tasks.
In the medical image domain, Huo et al. [30] introduced the Hierarchical Multiscale Feature Fusion Network (HiFuse), which combines global and local features using adaptive fusion modules to improve robustness in complex classification tasks. Chen et al. [31] developed a leukocyte detection model based on deformable-DETR with multi-level feature fusion, enhancing detection accuracy through detailed cellular characteristics. Lei et al. [32] presented the ConDSeg framework, employing a contrast-driven feature enhancement strategy for better boundary recognition in image segmentation, achieving state-of-the-art performance.
@@ 3.1 Overall Architecture
Fig. 1 illustrates the overall architecture of the CMSwinKAN model (micro version), which is designed to address the challenges of multi-scale feature representation and spatial dependency modeling in pathological image classification. This model is specifically tailored for the patch-level classification of pathological images from H&E-stained slides by combining a lightweight architecture with a strong interpretability.
The introduction of KAN layers significantly reduces model parameters by utilizing smaller embedding dimensions while enhancing feature representation and generalization [33]. The Swin KANsformer Block alternates between Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA) layers, replacing all multi-layer perception (MLP) layers with KAN layers, thereby boosting efficiency and interpretability.
The core structure of Contrastive-Driven Feature Augmentation (CDFA) in the CMSA module enhances multi-scale feature fusion by comparing attention across different scales, resulting in richer and more diagnostically meaningful insights [32]. These two modules work synergistically, enabling the CMSwinKAN model to efficiently and accurately classify H&E-stained pathological image patches and laying the foundation for WSI classification tasks.
@@ 3.2 Swin KANsformer Block
The Swin Transformer Block excels in image processing and medical image analysis compared to traditional vision Transformers due to its hierarchical feature maps and local window-based self-attention, which lower computational complexity and facilitate multi-scale feature extraction for high-resolution images. Its shifted window mechanism improves information exchange, but it has limitations like constrained global receptive fields and high computational demands. We address these issues with the Swin KANsformer Block, an optimized version of the original.
The Swin KANsformer Block replaces traditional MLP layers in the Swin Transformer Block with KAN layers[33], improving the model’s ability to process complex H&E-stained pathological images. The model divides input RGB H&E-stained images into non-overlapping 4×4 patches (48 features each) via the Patch Partition module before projecting them into a feature dimension $C$ using a linear embedding layer.
@@ 3.3 Contrastive-Driven Multi-Scale Aggregation Module
We analyze images at multiple scales to maintain spatial consistency. Our CMSA module effectively utilizes multi-scale features, incorporating the CDFA module to tackle challenges in medical image analysis.
@@
$\hat{\bm{V}}_{\Delta_{i,j}}=Softmax(\hat{A}_{1_{i,j}})\otimes(Softmax(\hat{A}_{2_{i,j}})\otimes(\bm{V}_{\Delta_{i,j}})),$ (3)
@@ 3.4 KAN Classification Head
The multi-scale features from the CMSA module are combined with output residuals from the Swin KANsformer’s 4th stage, maintaining detail and semantic information for better image classification. The KAN serves as the classification head to improve non-linear fitting and performance, with each activation function as a learnable spline:
@@
KAN uses B-spline basis functions to create piecewise polynomials with local support, enabling efficient handling of high-dimensional data. The B-spline basis function serves as the core component of KAN for constructing piecewise polynomials. For input data $x$ and a given set of grid points $t=\{t_{0},t_{1},\ldots,t_{n}\}$, the B-spline basis function $N_{i,k}\left(x\right)$ is computed as follows:
@@
For WSI-level pathological diagnosis, a heuristic soft voting classification process is developed that combines individual patch classifications with global feature information for a final label, using weighted slice importance for more robust results than hard voting, illustrated in Fig. 2. We also train an SVM model to identify specific tissue structures like neuropil and Schwannian stroma. Further details on the voting mechanism and SVM training will follow.
Figure 2: The architecture of the proposed WSI voting mechanism. After feature classification of the slices using an SVM classifier, weights are dynamically assigned based on classification confidence, ultimately implementing a weighted soft voting strategy for decision.